OpenStudy (anonymous):

who knows the equation of a parabola with a vertex (0,5) and a directrix=8

OpenStudy (anonymous):

Find the equation of the parabola given vertex at (5, 0) and y = -8 as dirctrix. The axis of symmetry is parallel to the Y –axis. The equation for the directrix is y = k – p. Substituting k = 0 and y = -8 to the equation : -8 = 0 – p -8 = -p 8 = p or p = 8 Therefore the equation of the parabola we are looking for is : (x – 5 ) ^ 2 = 8y

OpenStudy (anonymous):

thanks but the vertex was at (0,5) not (5,)

OpenStudy (anonymous):

not (5,0)

OpenStudy (anonymous):

Since the directrix is vertical, that means this is a sideways parabola (the quadratic term is y)

OpenStudy (anonymous):

the equation should be 4p(x-h) = (y-k)^2

OpenStudy (anonymous):

p=3 h=0 k=5 so the equation is 12x = (y-5)^2

OpenStudy (anonymous):

okay thanks. i have four answers they are: x=32(y-5)^2, y=1/32(x+8)^2, x=-1/32(y-5)^2, or

OpenStudy (anonymous):

this is on a multiple choice hw

OpenStudy (anonymous):

sorry p=8 so the answer is 32x=(y-5)^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sorry again... p=-8 so the equation is -32x=(y-5)^2 choose the equivalent equation in your multiple choice,...

OpenStudy (alexwee123):

is directrix at x=8 or y=8?

OpenStudy (anonymous):

haha... i didn't even see that "directrix=8"

OpenStudy (anonymous):

I assumed x=8